The faculty is getting ready for the senior/faculty basketball game, they have 20 interested staff. How many different ways can they pick a forward, center and guard?
Answer by Theo(13342)

20p3 = n! / (n-x)! = 20! / 17! = (20 * 19 * 18) * 17! / 17! = 20 * 19 * 18 = 6840

look at it another way.

there are 20 ways to pick a forward.
once the forward is picked there are 19 ways to pick a center.
once the center is picked there are 18 ways to pick a guard.

the number of possible ways is therefore 20 * 19 * 18 = 6840.